OpenStudy (anonymous):

solve for x: -2(x+3)=-2(x+1)-4

OpenStudy (anonymous):

@johnN.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@jeweaver145

OpenStudy (anonymous):

ok same process as the last problem, distributing the -2 into the parentheses and then combining like terms and then solving for x.

OpenStudy (anonymous):

hmm actually, are you sure you wrote that out correctly? it simplifies to 0=0..

OpenStudy (johnn.):

Sorry, I was offline. Let's solve your equation step-by-step. −2(x+3)=−2(x+1)−4 Step 1: Simplify both sides of the equation. −2(x+3)=−2(x+1)−4 Simplify: −2x−6=−2x−6 Step 2: Add 2x to both sides. −2x−6+2x=−2x−6+2x −6=−6 Step 3: Add 6 to both sides. −6+6=−6+6 0=0